# Composing Distributed Computations Through Task and Kernel Fusion

### v1

@@ 6.4. Qualitative Benefits
We note several qualitative benefits of our system architecture in contrast to approaches that attempt to optimize distributed programs entirely through analysis of imperative code. A key design decision of Diffuse is to leverage a distributed data model in a scale-free IR of computation that enables cheap dependence analysis between distributed computations. Separating out the reasoning about distributed computation avoids intertwining loop optimizations with distributed communication analyses. This separation also allows for information gained during the distributed analysis phase to be used in code generation: properties such as array non-aliasing are provided to the MLIR optimization passes to generate better code. Finally, the separation of distributed computation into tasks means that Diffuse does not need to identify optimizable fragments of static source code.
@@ 7. Evaluation
Overview. We evaluate Diffuse on unmodified, open source cuNumeric and Legate Sparse applications (links removed for double blind review), from micro-benchmarks to full applications. Many of these applications have appeared in prior publications (Bauer and Garland, 2019; Yadav et al., 2023), and range from tens to thousands of lines of Python. An overview of each application is in Figure 13. We compare each application’s performance when run with and without Diffuse. For some applications, we compare against manually optimized implementations by the original authors, and against the industry-standard PETSc (Balay et al., 2022) library for distributed sparse linear algebra. We show that when fusion opportunities are available, Diffuse can exploit them to find speedups in unmodified, distributed applications. Diffuse enables high-level programs to equal, and in many cases improve on, the performance of hand-optimized code.
We do not ablate on the optimizations in Section 5, as temporary elimination is essential for speedup with kernel fusion and memoization is a requirement for a practical implementation. We do not compare against the work of Sundram et al. (Sundram et al., 2022), which performs only task-fusion, as the version of cuNumeric they used is older and would not be a fair comparison. However, we have evaluated Diffuse with only task fusion and found that it did not yield speedups on our benchmarks. Task fusion alone can only reduce runtime overhead, and the task granularity of our benchmarks is larger than the minimum effective task granularity (Slaughter et al., 2020) of Legion. Finally, our benchmarks issue index tasks that have one point per GPU, so computations are not over-decomposed.
Benchmark Tasks/Iteration Tasks/Iteration (Fused) Window Size
Black-Scholes 67 1 70
Jacobi 3 2 5
CG 12.1 4.1 10
CGS 21.1 6.1 15
BiCGSTAB 27.1 8.1 15
GMG 24.1 11.1 15
CFD 378 40.7 30
TorchSWE 276.5 152.8 30
Figure 13. Index tasks executed per iteration with and without fusion. Task count is not always whole as iterations may launch different tasks, or fusion occurs across iteration boundaries. Window size was selected by Diffuse.
@@
(c) CG
(d) BiCGSTAB
(e) CGS
(f) GMG
(g) CFD
(h) TorchSWE
Figure 14. Weak scaling results across all applications (higher is better).
7.1. Weak Scaling Experiments
Black-Scholes. Black-Scholes is a trivially parallel benchmark that is a reference point on potential improvement when the entire application is fusible. Figure 14(a) shows that Diffuse achieves a 10.7x speedup over the unfused program on 128 GPUs, as the fused program is a single task containing a single GPU kernel making one pass over the data, greatly increasing the arithmetic intensity of the computation.
@@ 7.1. Weak Scaling Experiments
Sparse Krylov Solvers. We evaluate sparse Krylov solvers implemented with cuNumeric and Legate Sparse, namely Conjugate Gradient (CG), Conjugate Gradient Squared (CGS) and Bi-Conjugate Gradient Stabilized (BiCGSTAB). To perform a controlled comparison against PETSc, we modify Legate Sparse to perform a similar optimization as PETSc, where the non-zero coordinates in each sparse matrix partition are stored as 32-bit integers instead of 64-bit integers. [^1] PETSc stores coordinates in 32-bit integers even when 64-bit integers are requested at build time, affecting the performance of the SpMV kernel.
@@
The original implementation of CG in Legate Sparse had been optimized manually to perform many of the optimizations that Diffuse does automatically. As a result, the implementation no longer resembled the high-level description of CG. We compare against this manually fused implementation, a naturally written implementation using cuNumeric and Legate Sparse, and PETSc. Figure 14(c) shows that Diffuse automatically optimizes the naturally written CG so that it runs faster than both the manually optimized version and PETSc. Diffuse finds additional fusion opportunities by fusing AXPY’s and dot-products from different iterations.
@@
We implement unfused versions of CGS and BiCGSTAB in cuNumeric and Legate Sparse and compare to PETSc. Figures 14(e) and 14(d) show that Diffuse accelerates the high-level implementations of CGS and BiCGSTAB to outperform the unfused versions by 1.49x and 1.31x on average (geo-mean) and PETSc by 2.28x and 1.15x on average (geo-mean). PETSc exposes several fused kernels to users for use in building iterative solvers, but these kernels can quickly become complicated and esoteric [^2] Such as VecAXPBYPCZ in BiCGSTAB (https://petsc.org/main/manualpages/Vec/VecAXPBYPCZ/).. In contrast, Diffuse enables users to write high-level programs in cuNumeric and Legate Sparse and then derives optimized kernels for efficient execution.
@@
Geometric Multi-Grid Solver (GMG). Moving from smaller benchmarks to full applications, we apply Diffuse to a Geometric Multi-Grid (GMG) solver developed in Legate Sparse. The GMG solver is a CG-based iterative solver with a V-cycling preconditioner, the injection restriction operator, and a weighted Jacobi smoother. As with the previous benchmarks, using Diffuse with the more complex solver required no changes to user-facing code, and results in a 1.2x speedup over the original implementation, as seen in Figure 14(f).
@@
Computational Fluid Dynamics (CFD). We apply Diffuse to a cuNumeric application that solves the Navier-Stokes equations for 2D channel flow (Barba and Forsyth, 2019). The application performs element-wise operations on aliasing slices of distributed arrays, exposing opportunities for fusion. Diffuse finds between 1.8x–2.3x speedup over the original implementation, as shown in Figure 14(g). Diffuse achieves higher speedup on a single GPU than on multiple GPUs. On a single GPU, data is not partitioned, enabling longer sequences of tasks to satisfy fusion constraints. On multiple GPUs, the dependencies caused by aliasing data reduce the opportunities for fusion.
@@
Shallow Water Equation Solver (TorchSWE). Our final benchmark application is also our most complex: the cuNumeric port of the TorchSWE shallow-water equation solver (Chuang, 2021). We compare against the original cuNumeric port, as well as a version that the cuNumeric developers manually optimized using numpy.vectorize. The vectorize utility JIT-compiles a user-defined element-wise operator, doing some of the optimizations that Diffuse performs automatically. Figure 14(h) shows the performance of TorchSWE with Diffuse compared to these baselines. Diffuse achieves a 1.61x speedup on average (geo-mean) over the unfused TorchSWE, and a 1.35x speedup on average (geo-mean) over the manually vectorized version. Since Diffuse is analyzing the entire application, it can find fusion opportunities missed by developers optimizing the program by hand.
@@ 7.2. Compilation Time
We measure the overhead that Diffuse’s compilation imposes on overall runtime. When evaluating our benchmarks, we compute the throughput after warmup iterations have concluded. To measure the effect of compilation, we measure the warmup time with and without compilation. We then compute the number of iterations required for the fused version to be faster than the unfused version of the application when including the warmup compilation time. The results are shown in Figure 15; Diffuse’s compilation times are modest, requiring 25–119 iterations to amortize the cost of compilation. The fused Black-Scholes computation is so much faster than the unfused version that a single iteration is sufficient to amortize compilation. For Jacobi, compilation time was overlapped with expensive dense matrix-vector multiply kernel, and thus not exposed in the warmup. As seen in Figure 14(b), due to experimental variation, the fused and unfused versions of Jacobi are slightly faster or slower than each other on different GPU counts.
@@ 8. Related Work
Efficient Composition of Parallel Software. Diffuse aims to efficiently compose operations within and across distributed libraries. Some recent projects have tackled the problem of efficient composition; we discuss each in turn. Weld (Palkar et al., 2017) provides a loop-based IR in which users can define single-node library computations, and a runtime system that optimizes the IR to enable cross-function and cross-library optimizations. Split Annotations (Palkar and Zaharia, 2019) provides partitioning annotations for users to attach to library functions, and uses these annotations to run cache-sized batches of the functions to maximize data reuse. Both Weld and Split Annotations target a similar problem as Diffuse, but would require a model of distributed data like the one we propose to safely perform optimizations in a distributed setting. DaCe (Ben-Nun et al., 2019) is a compiler that leverages an IR called Stateful Dataflow MultiGraphs to perform optimizations like fusion on Python/NumPy programs. Distributed programs in DaCe are explicitly parallel, including manual communication with libraries like MPI, which requires different kinds of analyses. Jax (Bradbury et al., 2018) is a runtime system and compiler for NumPy-like operations that leverages the semantics of NumPy operations to statically perform transformations like fusion and automatic differentiation. In contrast to Jax, Diffuse is not tied to the semantics of any particular library, and leverages dynamic analyses to handle data-dependent control flow.
Distributed Runtime Systems. Diffuse uses a scale-free IR to efficiently perform distributed dependence and alias analyses. This is similar to Index Launches (Soi et al., 2021), a representation of distributed tasks that compresses the degree of parallelism. Diffuse’s model of distributed data supports content-based coherence, meaning that the same data may be referred to in multiple different ways. Legion (Bauer et al., 2012), which Diffuse builds upon, is a system that supports content-based coherence of distributed data. Legion exposes a more general interface for partitioning data, allowing a partition to contain arbitrary subsets. Legion then uses sophisticated algorithms for computing dependencies between tasks and maintaining coherence of distributed data (Bauer et al., 2023). Diffuse’s restricted representation and goal of fusion enable compact analyses for the dependence and coherence problems. In systems without content-based coherence, simpler approaches than ours may suffice, as aliasing distributed data is no longer a concern.
@@ 9. Conclusion
We introduced Diffuse, a system that performs task and kernel fusion on streams of distributed tasks, enabling optimizations that improve data reuse and remove allocations of distributed data structures in end user programs. Diffuse leverages a scale-free intermediate representation of distributed computation and data to perform these analyses in a scalable manner. These techniques enable Diffuse to compose computations in and across cuNumeric and Legate Sparse, matching or exceeding the performance of hand-tuned code.

### v2

@@ 6.4. Qualitative Benefits
We note several qualitative benefits of our system architecture in contrast to approaches that attempt to optimize distributed programs entirely through analysis of imperative code. A key design decision of Diffuse is to leverage a distributed data model in a scale-free IR of computation that enables cheap dependence analysis between distributed computations. Separating out the reasoning about distributed computation avoids intertwining loop optimizations with distributed communication analyses, allowing the loop optimizations to remain unaware of the distributed context. This separation also allows for information gained during the distributed analysis phase to be used in code generation: properties such as array non-aliasing are provided to the MLIR optimization passes to generate better code. Finally, the separation of distributed computation into tasks means that Diffuse does not need to identify optimizable fragments of static source code.
@@ 7. Evaluation
Overview. We evaluate Diffuse on unmodified, open source cuPyNumeric and Legate Sparse applications, from microbenchmarks to full applications. Many of these applications have appeared in prior publications (Bauer and Garland, 2019; Yadav et al., 2023), and range from tens to thousands of lines of Python. The unique capabilities of cuPyNumeric and Legate Sparse enable these pure Python applications with dynamic and data-dependent behavior to be scaled across multiple nodes of multiple GPUs. An overview of each application is in Figure 13. We compare each application’s performance when run with and without Diffuse — no changes to the application are needed to enable Diffuse. For some applications, a suitable baseline written in the industry-standard PETSc (Balay et al., 2022) library for distributed sparse linear algebra already exists, and we compare against those baselines. For other applications, we compare against manually optimized implementations by the original authors. However, some full cuPyNumeric applications have no baseline other than when run without name—these applications are sufficiently complex that developing an independent high-performance distributed, multi-GPU implementation is not feasible. We show that when fusion opportunities are available, Diffuse can exploit them to find speedups in unmodified, distributed applications. Diffuse enables high-level programs to equal, and in many cases improve on, the performance of hand-optimized code.
We do not ablate on the optimizations in Section 5, as temporary elimination is essential for speedup with kernel fusion and memoization is a requirement for a practical implementation. We do not compare against the work of Sundram et al. (Sundram et al., 2022), which performs only task-fusion, as the version of cuPyNumeric they used is older and would not be a fair comparison. However, we have evaluated Diffuse with only task fusion and found that it did not yield speedups on our benchmarks. Task fusion alone can only reduce runtime overhead, and the task granularity of our benchmarks is larger than the minimum effective task granularity (Slaughter et al., 2020) of Legion (1ms per task). The window sizes shown in Figure 13 were selected automatically by Diffuse through a process that increases the window size when all tasks in the current window size were fused. As a result, these window sizes enable the maximum amount of fusion possible in each application. Finally, our benchmarks issue index tasks that have one point per GPU, so computations are not over-decomposed.
Benchmark Tasks per Iteration Tasks per Iteration (Fused) Avg Task Length (ms) Window Size
Black-Scholes 67 1 5.3 70
Jacobi 3 2 5.3 5
CG 12.1 4.1 1.9 10
BiCGSTAB 27.1 8.1 1.7 15
GMG 24.1 11.1 1.8 15
CFD 378 40.7 1.1 30
TorchSWE 276.5 152.8 1.4 30
Figure 13. Index tasks per iteration with and without fusion. Task count is not always whole as iterations may launch different tasks, or fusion occurs across iteration boundaries. Reported task granularities are from unfused single-GPU executions. Window size was selected by Diffuse.
7.1. Weak Scaling Experiments
Black-Scholes. The Black-Scholes option pricing benchmark is a trivially-parallel micro-benchmark that contains a sequence of 67 data-parallel, and thus fusible, operations. It is a micro-benchmark that provides a reference point on potential improvement when the entire application is amenable to fusion. Figure 14(a) shows that Diffuse achieves a 10.7x speedup over the unfused program on 128 GPUs, as the fused program is a single task containing a single GPU kernel making one pass over the data, greatly increasing the arithmetic intensity of the computation.
@@ 7.1. Weak Scaling Experiments
Figure 14. Microbenchmark weak scaling (higher is better).
(a) CG
(b) BiCGSTAB
Figure 15. Weak scaling of linear solvers (higher is better).
(a) GMG
(b) CFD
(c) TorchSWE
Figure 16. Weak scaling of full applications (higher is better).
@@
Sparse Krylov Solvers. We evaluate sparse Krylov solvers implemented with cuPyNumeric and Legate Sparse, namely Conjugate Gradient (CG) and Bi-Conjugate Gradient Stabilized (BiCGSTAB). The PETSc benchmark implementations are implemented in MPI+C using PETSc’s API. To perform a controlled comparison against PETSc, we modify Legate Sparse to perform a similar optimization as PETSc, where the non-zero coordinates in each sparse matrix partition are stored as 32-bit integers instead of 64-bit integers. [^1] PETSc stores coordinates in 32-bit integers even when 64-bit integers are requested at build time, affecting the performance of the SpMV kernel.
@@
The original implementation of CG in Legate Sparse had been optimized manually to perform many of the optimizations that Diffuse does automatically. As a result, the implementation no longer resembled the high-level description of CG. We compare against this manually fused implementation, a naturally written implementation using cuPyNumeric and Legate Sparse, and PETSc. Figure 15(a) shows that Diffuse automatically optimizes the naturally written CG so that it runs faster than both the manually optimized version and PETSc. Diffuse finds additional fusion opportunities by fusing AXPY’s and dot-products from different iterations.
@@
We implement an unfused version of BiCGSTAB in cuPyNumeric and Legate Sparse and compare against PETSc. Figure 15(b) shows that Diffuse accelerates the high-level implementation of BiCGSTAB to outperform the unfused version by 1.31x on average (geo-mean) and PETSc by 1.15x on average (geo-mean). PETSc exposes several fused kernels to users for use in building iterative solvers, but these kernels can quickly become complicated and esoteric [^2] Such as VecAXPBYPCZ in BiCGSTAB (https://petsc.org/main/manualpages/Vec/VecAXPBYPCZ/).. In contrast, Diffuse enables users to write high-level programs in cuPyNumeric and Legate Sparse and then derives optimized kernels for efficient execution.
@@
Geometric Multi-Grid Solver (GMG). Moving from smaller benchmarks to full applications, we apply Diffuse to a Geometric Multi-Grid (GMG) solver developed in Legate Sparse. The GMG solver is a CG-based iterative solver with a V-cycling preconditioner, the injection restriction operator, and a weighted Jacobi smoother. As with the previous benchmarks, using Diffuse with the more complex solver required no changes to user-facing code, and results in a 1.2x speedup over the original implementation, as seen in Figure 16(a).
@@
Computational Fluid Dynamics (CFD). We apply Diffuse to a cuPyNumeric application that solves the Navier-Stokes equations for 2D channel flow (Barba and Forsyth, 2019). The application performs element-wise operations on aliasing slices of distributed arrays, exposing opportunities for fusion. Diffuse finds between 1.8x–2.3x speedup over the original implementation, as shown in Figure 16(b). Diffuse achieves higher speedup on a single GPU than on multiple GPUs. On a single GPU, data is not partitioned, enabling longer sequences of tasks to satisfy fusion constraints. On multiple GPUs, the dependencies caused by aliasing data reduce the opportunities for fusion.
@@
Shallow Water Equation Solver (TorchSWE). Our final benchmark application is also our most complex: the cuPyNumeric port of the TorchSWE shallow-water equation solver (Chuang, 2021). We compare against the original cuPyNumeric port, as well as a version that the cuPyNumeric developers manually optimized using numpy.vectorize. The vectorize utility JIT-compiles a user-defined element-wise operator, doing some of the optimizations that Diffuse performs automatically. Figure 16(c) shows the performance of TorchSWE with Diffuse compared to these baselines. Diffuse achieves a 1.61x speedup on average (geo-mean) over the unfused TorchSWE, and a 1.35x speedup on average (geo-mean) over the manually vectorized version (labeled with “Manually Fused” in Figure 16(c)). Since Diffuse is analyzing the entire application, it can find fusion opportunities missed by developers optimizing the program by hand.
@@ 7.2. Compilation Time
We measure the overhead that Diffuse’s compilation imposes on overall runtime. When evaluating our benchmarks, we compute the throughput after warmup iterations have concluded. To measure the effect of compilation, we measure the warmup time with and without compilation, using the window sizes reported in Figure 13. We then compute the number of iterations required for the fused version to be faster than the unfused version of the application when including the warmup compilation time. The results are shown in Figure 17; Diffuse’s compilation times are modest, requiring 25–119 iterations to amortize the cost of compilation. The fused Black-Scholes computation is so much faster than the unfused version that a single iteration is sufficient to amortize compilation. For Jacobi, compilation time was overlapped with expensive dense matrix-vector multiply kernel, and thus not exposed in the warmup. As seen in Figure 14(b), due to experimental variation, the fused and unfused versions of Jacobi are slightly faster or slower than each other on different GPU counts. These costs are especially reasonable as scientific applications like the ones we evaluated would be run in production for thousands to millions of iterations. In the future, a production-grade implementation of Diffuse could maintain a cache of compiled kernels on disk, rather than in memory, and pay the compilation cost only the first time the application is run.
@@ 8. Related Work
Efficient Composition of Parallel Software. Diffuse aims to efficiently compose operations within and across distributed libraries. Some recent projects have tackled the problem of efficient composition; we discuss each in turn. Weld (Palkar et al., 2017) provides a loop-based IR in which users can define single-node library computations, and a runtime system that optimizes the IR to enable cross-function and cross-library optimizations. Split Annotations (Palkar and Zaharia, 2019) provides partitioning annotations for users to attach to library functions, and uses these annotations to run cache-sized batches of the functions to maximize data reuse. Both Weld and Split Annotations target a similar problem as Diffuse, but would require a model of distributed data like the one we propose to safely perform optimizations in a distributed setting. DaCe (Ben-Nun et al., 2019) is a compiler that leverages an IR called Stateful Dataflow MultiGraphs to perform optimizations like fusion on Python/NumPy programs. Distributed programs in DaCe are explicitly parallel, including manual communication with libraries like MPI, which requires different kinds of analyses.
Jax (Bradbury et al., 2018) and PyTorch (Ansel et al., 2024) are machine-learning systems that compile NumPy-like descriptions of neural networks to perform optimizations like fusion and automatic differentiation. Systems like Jax and PyTorch accept structured program representations (neural network graphs) and apply optimizations that leverage domain-specific knowledge, many of which are not possible for Diffuse to perform. In contrast, Diffuse only leverages the privilege information about tasks to perform optimizations, and allows for description of programs with complex aliasing and mutation that are not possible to represent in ML systems, like the CFD or TorchSWE simulations. We consider Diffuse to be a different point in the design space than these ML systems, focusing on fusion in a more general setting without application-specific knowledge.
Distributed Runtime Systems. Diffuse uses a scale-free IR to efficiently perform distributed dependence and alias analyses. This is similar to Index Launches (Soi et al., 2021), a representation of distributed tasks that compresses the degree of parallelism. Diffuse’s model of distributed data supports content-based coherence, meaning that the same data may be referred to in multiple different ways. Legion (Bauer et al., 2012), which Diffuse builds upon, is a system that supports content-based coherence of distributed data. Legion exposes a more general interface for partitioning data, allowing a partition to contain arbitrary subsets. Legion then uses sophisticated algorithms for computing dependencies between tasks and maintaining coherence of distributed data (Bauer et al., 2023). Legion’s flexible data model and support for precise dependence analysis at scale are critical features for building libraries like cuPyNumeric and Legate Sparse. Supporting Legion’s flexible data model is a key challenge in Diffuse, as libraries that target Legion depend on this capability. Diffuse’s restricted data representation and goal of only fusion enable compact analyses for the dependence and coherence problems. In systems without content-based coherence, simpler approaches than ours may suffice, as aliasing distributed data is no longer a concern.
@@ 9. Conclusion
We introduced Diffuse, a system that performs task and kernel fusion on streams of distributed tasks, enabling optimizations that improve data reuse and remove allocations of distributed data structures in end user programs. Diffuse leverages a scale-free intermediate representation of distributed computation and data to perform these analyses in a scalable manner. These techniques enable Diffuse to compose computations in and across cuPyNumeric and Legate Sparse, matching or exceeding the performance of hand-tuned code.
Acknowledgements
We thank Scott Kovach for his assistance in formalizing the fusion correctness proof. We thank Shriram Jagannathan and Irina Demeshko for their assistance in running the vectorized TorchSWE benchmark. We thank (in no particular order) David Broman, James Dong, AJ Root, Scott Kovach, Parthiv Krishna, Benjamin Driscoll, Olivia Hsu, Marco Siracusa, Rubens Lacouture for their comments and discussions on early stages of this manuscript. Rohan Yadav was supported by an NVIDIA Graduate Fellowship, and part of this work was done while Rohan Yadav was an intern at NVIDIA Research. This work was in part supported by the National Science Foundation under Grant CCF-2216964.